# ScribblePrompt: Fast and Flexible Interactive Segmentation for Any Biomedical Image

### v2

@@ 5.2 Simulated Interactions
Results. Fig. 6 shows that both versions of ScribblePrompt outperform baseline methods for all simulated interaction procedures at all numbers of interactions. We show examples with simulated interactions in Fig. 1 and Fig. 5 (bottom). We show results with similar trends using bounding boxes and by dataset with fully-supervised baselines in Appendix F.
Discussion. Comparing the simulated scribble and click results (Fig. 6) highlights the efficiency of scribble-based interactions. Both ScribblePrompt models reach an average Dice above $0.8$ in two scribble steps. It takes five steps of carefully-placed center clicks or eight random clicks to reach the same average Dice score with ScribblePrompt-SAM.
@@
Figure 6: Simulated clicks and scribbles. We simulate interactions following three scribble protocols and three click protocols. We show more results, with similar trends, in Sec. F.2 (Fig. 21).
@@ 5.3 User Study
Table 2: User study. We report means and standard deviations for Dice score, HD95 and total time per segmentation, as well as the median (and max) number of times the participants ($n=16$) refreshed the prediction.
SAM (ViT-b) ScribblePrompt-UNet
Dice Score ($\uparrow$) $0.73\pm 0.30$ $0.84\pm 0.13$
HD95 ($\downarrow$) $16.1\pm 37.1$ $4.5\pm 9.1$
Sec./task ($\downarrow$) $131.7\pm 168.5$ $94.2\pm 89.0$
Iter./task ($\downarrow$) $6~{}(47)$ $3~{}(19)$
@@
Setup. Study participants ($n=16$) were neuroimaging researchers at an academic hospital. They used each model to segment a series of nine test images from different evaluation datasets. For each segmentation task, the participants were shown the target segmentation and were asked to prompt the model until the predicted segmentation closely matched the target, or they could no longer improve the prediction. We provided participants with the target segmentation to disentangle the cognitive process of identifying the region of interest from prompting the model to achieve the desired segmentation. We provide additional details and visualizations in Appendix G.
@@
Discussion. The most common factors that influenced participant preference for ScribblePrompt were 1) being able get accurate predictions from multiple types of prompts, including scribbles and 2) responsiveness to their corrections, which enabled fine-grained control over the next prediction. For some tasks, such as retinal vein segmentation, SAM was unable to make accurate predictions even with many corrections (Fig. 35). As a result, there was more variability in Dice score and time per segmentation, for SAM than ScribblePrompt.
@@ 5.4 Inference Runtime
Results. On a single CPU, ScribblePrompt-UNet requires $0.27\pm 0.04$ sec per prediction, enabling the model to be used in real-time even in low-resource environments. The next most accurate model (Fig. 6), SAM ViT-h, requires over 2 minutes per prediction on a CPU ($130.79\pm 7.96$ s). SAM ViT-b takes $13.59\pm 0.77$ seconds per prediction. We report results for all baselines and on GPU hardware with similar trends in Appendix H.
@@
Figure 7: Ablations. We report Dice after five steps of simulated interactions following three inference-time interaction procedures. Error bars show 95% CI. (a) shows mean change in Dice relative to ScribblePrompt-UNet trained without any synthetic labels ($p_{synth}=0$).
6 Ablations
@@ 6.1 Synthetic Labels
Setup. We trained ScribblePrompt-UNet with different values of $p_{synth}$, the probability of sampling a synthetic label.
Results. Fig. 7(a) shows the effect of varying the percentage of synthetic labels during training for ScribblePrompt-UNet. Training with both real and synthetic labels improves generalization to new datasets, compared to training with only real labels. Using $p_{synth}=0.5$ results in the highest Dice. Training with only synthetic labels results in worse Dice scores. We show similar results for ScribblePrompt-SAM and other interaction procedures in Sec. I.1.
@@ 6.2 Prompt Types
Results. Fig. 7(b) shows ScribblePrompt-UNet trained with scribbles, clicks and bounding boxes predicts segmentations more accurately than do ablated versions of ScribblePrompt-UNet. We show results for other interaction procedures with similar trends in Sec. I.2.
@@ 8 Acknowledgements
Appendix
Appendix A: Demo and Code
Appendix B: ScribblePrompt Implementation
Appendix D: Experimental Setup
Appendix C: Data
Appendix E: Evaluation with Manual Scribbles
Appendix F: Evaluation with Simulated Interactions
Appendix G: User Study
Appendix H: Inference Runtime
Appendix I: Ablations
Appendix A Demo and Code
@@
Appendix B ScribblePrompt Implementation
B.1 Prompt Simulation
@@ B.3 Synthetic Labels
Appendix C Data
We build on large dataset gathering efforts like MegaMedical [20, 109] to compile a collection of 77 open-access biomedical imaging datasets for training and evaluation, covering over 54k scans, 16 image types, and 711 labels. We gathered datasets with a particular focus on Microscopy, X-Ray, and Ultrasound modalities, which were not as well represented in the original MegaMedical [20]. The full list of datasets is provided in Tab. 4, Tab. 5, and Tab. 6.
@@ Appendix C Data
Division of Datasets. The division of datasets and subjects for training, model selection, and evaluation is summarized in Tab. 3. The 77 datasets were divided into 65 training datasets (Tables 5 and 6), 12 evaluation datasets. Data from 9 (out of 12) of the evaluation datasets was used in model development for model selection, and final evaluation. The other 3 evaluation datasets were completely held-out from model development and only used in the final evaluation.
@@
Table 5: Train datasets (part 1 of 2). We assembled the following set of datasets to train ScribblePrompt. For the relative size of datasets, we have included the number of unique scans (subject and modality pairs) that each dataset has.
@@
Table 6: Train datasets (part 2 of 2). We assembled the following set of datasets to train ScribblePrompt. For the relative size of datasets, we have included the number of unique scans (subject and modality pairs) that each dataset has.
Dataset Name Description Scans Modalities
@@
Appendix D Experimental Setup
@@
Task Diversity. The final ScribblePrompt-UNet and ScribblePrompt-SAM models were trained with $p_{synth}=0.5$. Tab. 7 shows the data augmentations we used, similar to the in-task augmentations from [20].
@@
SAM Baselines. For baseline methods using the SAM architecture, we evaluate the models in both “single mask” and “multi-mask” mode. For each baseline method and interaction procedure, we selected the best performing mode based on the average Dice across the validation data and report final results on test data using that mode. In the results with simulated clicks and scribbles by dataset in Sec. F.2, we show results using both modes. For ScribblePrompt-SAM and SAM-Med2D with adapter layers, multi-mask mode resulted in the highest Dice. For SAM-Med2D without adapter layers, we found multi-mask mode led to higher Dice for scribble inputs while single-mask mode led to higher Dice with click inputs. For SAM (ViT-b and ViT-h) and MedSAM, single-mask mode resulted in the higher Dice on average.
Appendix E Manual Scribbles
@@
E.1 Setup
MedScribble Dataset. We collected a diverse dataset of manual scribble annotations, which is available at scribbleprompt.csail.mit.edu. The MedScribble dataset contains annotations from 3 annotators for 64 image-segmentation pairs. The examples were randomly selected from the validation split of 14 different datasets (7 training datasets and 7 validation datasets) [14, 130, 68, 69, 59, 58, 43, 88, 48, 72, 117, 1, 105, 49, 150, 151].
@@ E.2 Results
Appendix F Simulated Interactions
We present additional results from experiments with simulated interactions (Sec. 5.2).
F.1 Bounding Boxes
@@ F.2 Scribbles and Clicks
Appendix G User Study
We conducted a user study comparing ScribblePrompt-UNet to SAM (ViT-b). We provide additional details on the design and implementation of the user study.
@@ Appendix G User Study
Appendix H Inference Runtime
@@
Results. We show performance results in Tab. 8. On a single CPU, ScribblePrompt-UNet requires $0.27\pm 0.04$ sec per prediction, enabling the model to be used in real-time even in low-resource environments. ScribblePrompt-UNet is faster than the baseline methods on a GPU. Unlike the transformer-based models, the inference time for each prediction is does not depend on the number of of input interactions for ScribblePrompt-UNet.
@@
Architecture   Param.   CPU Runtime (sec)   GPU Runtime (ms)
SAM (ViT-h) 641M $130.79\pm 7.96$ $504.36\pm 57.72$
SAM (ViT-b) 94M $13.59\pm 0.77$ $133.85\pm 24.26$
SAM-Med2D w/ adapter 271M $1.23\pm 0.07$ $35.06\pm 12.88$
SAM-Med2D w.o. adapter 91M $0.63\pm 0.02$ $24.86\pm 9.56$
MIDeepSeg 3M $0.08\pm 0.02$ $65.75\pm 21.87$
ScribblePrompt-UNet 4M $0.27\pm 0.04$ $1.96\pm 0.20$
Appendix I Ablations

### v3

@@ 5.2 Simulated Interactions
5.2.1 Results.
Fig. 6 shows that both versions of ScribblePrompt outperform baseline methods for all simulated interaction procedures at all numbers of interactions. We show examples with simulated interactions in Fig. 1 and Fig. 5 (bottom). We show results with similar trends using bounding boxes and by dataset with fully-supervised baselines in Supplementary Material Sec. F.
5.2.2 Discussion.
Comparing the simulated scribble and click results (Fig. 6) highlights the efficiency of scribble-based interactions. Both ScribblePrompt models reach an average Dice above $0.8$ in two scribble steps. It takes five steps of carefully-placed center clicks or eight random clicks to reach the same average Dice score with ScribblePrompt-SAM.
@@
Figure 6: Simulated clicks and scribbles. We simulate interactions following three scribble protocols and three click protocols. We show more results and example predictions, with similar trends, in Supplementary Material Sec. 0.F.2.
@@ 5.3 User Study
Setup. Study participants ($n=16$) were neuroimaging researchers at an academic hospital. They used each model to segment a series of nine test images from different evaluation datasets. For each segmentation task, participants were shown the target segmentation and were asked to prompt the model until the predicted segmentation closely matched the target, or they could no longer improve the prediction. We provided participants with the target segmentation to disentangle the cognitive process of identifying the region of interest from prompting the model to achieve the desired segmentation. We provide additional details and visualizations in Supplementary Material Sec. 0.G.
@@
Discussion. The most common factors that influenced participant preference for ScribblePrompt were 1) being able get accurate predictions from multiple types of prompts, including scribbles and 2) responsiveness to their corrections, which enabled fine-grained control over the next prediction. For some tasks, such as retinal vein segmentation, SAM was unable to make accurate predictions even with many corrections. As a result, there was more variability in Dice score and time per segmentation, for SAM than ScribblePrompt.
Table 2: User study. Mean $\pm$ std. for Dice score, HD95 and total time per segmentation. Median (and max) number of times the participants refreshed the prediction.
Dice Score ($\uparrow$) HD95 ($\downarrow$) Sec./task ($\downarrow$) Iter./task ($\downarrow$)
SAM (ViT-b) $0.73\pm 0.30$ $16.1\pm 37.1$ $131.7\pm 168.5$ $6~{}(47)$
ScribblePrompt-UNet $0.84\pm 0.13$ $4.5\pm 9.1$ $94.2\pm 89.0$ $3~{}(19)$
@@ 5.4 Inference Runtime
Results. On a single CPU, ScribblePrompt-UNet requires $0.27\pm 0.04$ sec per prediction, enabling the model to be used even in low-resource environments. The next most accurate model (Fig. 6), SAM ViT-h, requires over 2 minutes per prediction on a CPU ($130.79\pm 7.96$ sec). SAM ViT-b takes $13.59\pm 0.77$ seconds per prediction. We report results for all baselines and on GPU hardware with similar trends in Supplementary Material Sec. 0.H.
6 Ablations
@@
Figure 7: Ablations. We report Dice after five steps of simulated interactions following three inference-time interaction procedures. Error bars show 95% CI. (a) shows mean change in Dice relative to ScribblePrompt-UNet trained without any synthetic labels.
@@ 6.1 Synthetic Labels
Fig. 7(a) shows the effect of varying the probability of sampling a synthetic label during training for ScribblePrompt-UNet. Training with both real and synthetic labels improves generalization to new datasets, compared to training with only real labels. Using $p_{synth}=0.5$ results in the highest Dice. Training with only synthetic labels results in worse Dice scores. We show similar results for ScribblePrompt-SAM and other interactions in Supplementary Material Sec. 0.I.1.
@@ 6.2 Prompt Types
Results. Fig. 7(b) shows ScribblePrompt-UNet trained with scribbles, clicks and bounding boxes predicts segmentations more accurately than do ablated versions of ScribblePrompt-UNet. We show results for other interaction procedures with similar trends in Supplementary Material Sec. 0.I.2.
@@ Acknowledgements
Appendix 0.A: Demo and Code
Appendix 0.B: ScribblePrompt Implementation
Appendix 0.C: Data
Appendix 0.D: Experimental Setup
Appendix 0.E: Evaluation with Manual Scribbles
Appendix 0.F: Evaluation with Simulated Interactions
Appendix 0.G: User Study
Appendix 0.H: Inference Runtime
Appendix 0.I: Ablations
Appendix 0.A Demo and Code
@@
Appendix 0.B ScribblePrompt Implementation
0.B.1 Prompt Simulation
@@ 0.B.3 Synthetic Labels
Appendix 0.C Data
We build on large dataset gathering efforts like MegaMedical [20, 115] to compile a collection of 77 open-access biomedical imaging datasets for training and evaluation, covering over 54k scans, 16 image types, and 711 labels. We gathered datasets with a particular focus on Microscopy, X-Ray, and Ultrasound modalities, which were not as well represented in the original MegaMedical [20]. The full list of datasets is provided in Tab. 4 and Tab. 5.
@@ Appendix 0.C Data
Division of Datasets. The division of datasets and subjects for training, model selection, and evaluation is summarized in Tab. 3. The 77 datasets were divided into 65 training datasets (Table 5, 12 evaluation datasets. Data from 9 (out of 12) of the evaluation datasets was used in model development for model selection, and final evaluation. The other 3 evaluation datasets were completely held-out from model development and only used in the final evaluation.
@@
Table 5: Training datasets. We assembled the following set of datasets to train ScribblePrompt. For the relative size of datasets, we have included the number of unique scans (subject and modality pairs) that each dataset has.
@@
Appendix 0.D Experimental Setup
@@
Task Diversity. The final ScribblePrompt-UNet and ScribblePrompt-SAM models were trained with $p_{synth}=0.5$. Tab. 6 shows the data augmentations we used, similar to the in-task augmentations from [20, 115].
@@
SAM Baselines. For baseline methods using the SAM architecture, we evaluate the models in both “single mask” and “multi-mask” mode. For each baseline method and interaction procedure, we selected the best performing mode based on the average Dice across the validation data and report final results on test data using that mode. In the results with simulated clicks and scribbles by dataset in Sec. 0.F.2, we show results using both modes. For ScribblePrompt-SAM and SAM-Med2D with adapter layers, multi-mask mode resulted in the highest Dice. For SAM-Med2D without adapter layers, we found multi-mask mode led to higher Dice for scribble inputs while single-mask mode led to higher Dice with click inputs. For SAM (ViT-b and ViT-h) and MedSAM, single-mask mode resulted in the higher Dice on average.
Appendix 0.E Manual Scribbles
@@
0.E.1 Setup
MedScribble Dataset. We collected a diverse dataset of manual scribble annotations, which is available at https://scribbleprompt.csail.mit.edu/data. The MedScribble dataset contains annotations from 3 annotators for 64 image segmentation pairs. The examples were randomly selected from the validation split of 14 different datasets (7 training datasets and 7 validation datasets) [14, 135, 69, 70, 59, 60, 45, 94, 50, 73, 123, 1, 111, 51, 156, 157].
@@ 0.E.2 Results
0.E.3 Comparison to Scribble-Supervised Learning
We report preliminary results comparing ScribblePrompt to scribble-supervised learning. Scribble-supervised learning methods use scribble annotations as supervision to train automatic segmentation models for predicting segmentation given only an input image [76, 74, 151, 86, 40]. These models are task-specific; a new model must be trained using scribble-supervised learning for each new task and training requires many scribble-annotated images from the same task to produce accurate results. In contrast, ScribblePrompt can perform new segmentation tasks at inference time without retraining, using scribbles as input.
Setup. We compare ScribblePrompt-UNet to ScribFormer [75], a recent state-of-the-art scribble-supervised learning method, on the ACDC scribbles dataset [14]. Experiments reported in [75] show that ScribFormer’s performance varies with the amount of training data, from 0.847 Dice given 14 training subjects to 0.894 Dice given 70 training subjects (and 15 validation subjects) from ACDC.
We evaluate each method given the same test data as in our manual scribbles evaluation: 20 subjects with scribble-annotations for three labels and background. For ScribFormer, we randomly partition the 20 test subjects into 80% train and 20% validation by subject, and train following [75]. We run inference for each model on all 20 test subjects, and report results averaged across the three labels for the 380 slices.
Results. Tab. 7 shows the difference in mean Dice between ScribblePrompt-UNet and ScribFormer is not statistically significant ($p=0.70$ with a paired t-test). Training ScribFormer required 2 hours using a Nvidia A100 GPU with 16 CPUs.
Table 7: Comparison to scribble-supervised learning. Mean Dice and HD95 with 95% CI of predicted segmentations for ACDC ($n=1,140$).
$\uparrow$ Dice Score $\downarrow$ HD95
ScribFormer $0.85\pm 0.01$ $4.05\pm 0.99$
ScribblePrompt-UNet $0.84\pm 0.01$ $1.80\pm 0.11$
Discussion. Given limited scribble-annotated data from ACDC, ScribblePrompt-UNet predicts segmentations with similar Dice scores and lower HD95 compared to a scribble-supervised learning model trained on the data.
Appendix 0.F Simulated Interactions
We present additional results from the experiments in Sec. 5.2 with simulated interactions.
0.F.1 Bounding Boxes
@@ 0.F.2 Scribbles and Clicks
Appendix 0.G User Study
We conducted a user study comparing ScribblePrompt-UNet to SAM (ViT-b). We provide additional details on the user study design and implementation.
@@ Appendix 0.G User Study
Appendix 0.H Inference Runtime
@@
Results. We show performance results in Tab. 8. On a single CPU, ScribblePrompt-UNet requires $0.27\pm 0.04$ sec per prediction, enabling the model to be used even in low-resource environments. Prior work on interactive interfaces indicates that $<0.5$ sec latency is sufficient for cognitive tasks [82]. ScribblePrompt-UNet is also faster than the baseline methods on a GPU.
@@
Architecture Param. CPU Runtime (sec) GPU Runtime (ms) GPU Memory
SAM (ViT-h) 641M $130.79\pm 7.96$ $504.36\pm 57.72$ $21.912$ GB
SAM (ViT-b) 94M $13.59\pm 0.77$ $133.85\pm 24.26$ $7.144$ GB
SAM-Med2D w/ adapter 271M $1.23\pm 0.07$ $35.06\pm 12.88$ $1.489$ GB
SAM-Med2D w.o. adapter 91M $0.63\pm 0.02$ $24.86\pm 9.56$ $734$ MB
MIDeepSeg 3M $0.08\pm 0.02$ $65.75\pm 21.87$ $11$ MB
ScribblePrompt-UNet 4M $0.27\pm 0.04$ $1.96\pm 0.20$ $125$ MB
Appendix 0.I Ablations